OpenStudy (anonymous):

Identify the solution set of the inequality using the given replacement set. x < –7; {–15, –10.5, –7.2, –7, 0, 6}

OpenStudy (anonymous):

0 and 6 should be replaced. For x<-7, and 0 and 6 are greater than -7.

OpenStudy (anonymous):

can u help me on like 3 more???

OpenStudy (anonymous):

I think that -7 should be replaced too, for it is equal to -7, not smaller than it. Maybe it is x≤−7 , not x<-7.

OpenStudy (anonymous):

Solve. 4y + 16 > –8

OpenStudy (anonymous):

First, let's work on the left hand side of your inequality, the 4y+16 This means, for instance, to see if it can be simplified at all. Multiply y and 4 Multiply y and 1 The y just gets copied along. The answer is y y 4*y evaluates to 4y 4*y+16 evaluates to 4y+16

OpenStudy (anonymous):

So, all-in-all, the left hand side of your inequality can be written as: 4y+16

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Now, let's work on the right hand side of your inequality, the -8 Because of the minus sign 8 becomes - 8 The answer is -8 The right hand side of your inequality can be written as: -8

OpenStudy (anonymous):

So with these (any) simplifications, the inequality we'll set out to solve is: 4y+16 › -8 Move the 16 to the right hand side by subtracting 16 from both sides, like this: From the left hand side: 16 - 16 = 0 The answer is 4y From the right hand side: -8 - 16 = -24 The answer is -24 Now, the inequality reads: 4y › -24 To isolate the y, we have to divide both sides of the inequality by the other "stuff" (variables or coefficients) around the y on the left side of the inequality.

OpenStudy (anonymous):

The last step is to divide both sides of the inequality by 4 like this: To divide y by 1 The y just gets copied along in the numerator. The answer is y 4y ÷ 4 = y -24 ÷ 4 = -6 The solution to your inequality is: y › -6

OpenStudy (anonymous):

solve -r/4 < 8

OpenStudy (anonymous):

First, let's work on the left hand side of your inequality, the -r/4 This means, for instance, to see if it can be simplified at all. To divide r by 1 The r just gets copied along in the numerator. The answer is r r/4 evaluates to Because of the minus sign becomes - The answer is So, all-in-all, the left hand side of your inequality can be written as: Now, let's work on the right hand side of your inequality, the 8 The right hand side of your inequality can be written as: 8 So with these (any) simplifications, the inequality we'll set out to solve is: ‹ 8 To isolate the r, we have to divide both sides of the inequality by the other "stuff" (variables or coefficients) around the r on the left side of the inequality. The last step is to divide both sides of the inequality by like this: To divide r by 1 The r just gets copied along in the numerator. The answer is r ÷ = r 8 ÷ = -32 Note! Here's an important point about inequalities: We are now dividing by -0.25, which is a number less than zero! This changes the form of your inequality. Or, in other words, your ‹ will now become a ›. The solution to your inequality is: r › -32 So, your solution is: r must be greater than -32

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

and last how would this look on a number line -x/2<8